# Conventional Quantum Theory Does Not Support A Coherent Relational Account

### v1

@@
ABSTRACT. I review a counterexample to the frequent claim that discrepancies among observers resulting from conventional quantum theory’s inability to define “measurement,” such as those arising in the Wigner’s Friend thought experiment, remain private and incommensurable. I consider the implications for a recent attempt to shield Relational Quantum Mechanics from such inconsistencies and conclude that it is not successful.
@@ 1. Introduction and Background
This paper reviews the challenge to conventional quantum theory (which I will henceforth call CQM) from inconsistencies arising from its prediction of macroscopic superpositions such as Schrödinger’s Cat (SC) and Wigner’s Friend (WF). I focus on an interpretational approach first proposed by Rovelli (1996) called “Relational Quantum Mechanics”(henceforth RQM). RQM is revisited with modifications in Adlam and Rovelli 2022 (henceforth AR2022), The latter paper is the subject of the present critique.
@@
The topic at hand is therefore fundamentally the measurement problem (MP) of CQM, since thought experiments involving ‘absurd’ macroscopic superpositions (the characterization in Frauchiger and Renner 2018, and clearly Schrödinger’s intent) are all illustrations of the MP. It is important to keep that point in mind in what follows. For clarity, CQM is the standard approach to quantum theory that denies any real physical non-unitarity and therefore must treat observed determinate outcomes as brute empirical facts with no physical explanation within the theory (more strongly, observed outcomes fail to be predicted by the theory, as the MP macroscopic superposition paradoxes illustrate). This point is granted by AR2022, who say (in the context of a WF experiment):
@@
“Unitary quantum mechanics does not provide any mechanism for a single measurement outcome to be selected and actualised for Alice in the first place, so it certainly cannot tell us anything about the relationship between her outcome and Bob’s outcome - this is a question which lies entirely outside the unitary part of the theory…” (AR2022)
@@
The locution ‘unitary part of the theory’ might seem to suggest that there is another ‘part of the theory,’ but that merely leaves room for an ad hoc application of a ‘collapse postulate’ notoriously subject to the ‘Shifty Split’ between the system to which CQM is taken as applying and some putative ‘outside observer’. This situation is what gives rise to the Wigner’s Friend and related inconsistencies such as that of the Frauchiger-Renner Paradox (Frauchiger and Renner, 2018). Of course, there are many attempts to deal with the MP and resulting inconsistencies arising from CQM and its Shifty Split. The paper critiqued herein, AR2022, is one of those attempts; I argue in what follows that it fails.
@@
In contrast to CQM, formulations involving objective reduction or collapse (henceforth OR) [^1] Some examples are the gravitational collapse model of Penrose 2000 and the Transactional Interpretation (Cramer 1986, most recent relativistic version in Kastner 2022) are not subject to the MP and thus do not lead to macroscopic superpositions such as that of Bob and his lab in the Wigner’s Friend paradox. This is the case regardless of one’s opinion of the theoretical or practical adequacy of such theories, which is a separate issue. Interestingly, one of the authors of AR2022 (EA) does acknowledge that a particular OR formulation solves the MP in a footnote to a paper entitled “Do We Have Any Viable Solution to the Measurement Problem?”, where she says:
@@
“…the transactional interpretation…is a solution to the measurement problem which has the property of being retrocausal.” (Adlam 2023). [^2] Adlam’s references are out of date, however; in addition to Cramer (1986), a more recent, fully relativistic account is Kastner (2022); see also Kastner (2018) in regards to TI’s solution to the MP.
@@
(This comment was made as an aside in the context of properties a particular formulation or interpretation might have. Oddly, however, although acknowledged by the author as a solution to the MP in the quoted footnote, the transactional interpretation is not mentioned anywhere in the body of the paper, despite its title—even with the emphasis on “Any”). [^3] This phenomenon, in which an author disregards a formulation that they themselves admit in a footnote solves the problem that is the subject of their paper, would provide for an interesting study in the sociology and psychology of science. However, such considerations are beyond the scope of this paper.
@@
Thus, at the risk of repetition (for some emphasis seems necessary), OR formulations eliminate the MP so that such ‘paradoxes’ as Schrödinger’s Cat and Wigner’s Friend do not arise in the first place. It is important to keep this point—that the notorious observer-related paradoxes do not arise under OR formulations—firmly in mind, since it appears to have been somehow forgotten in AR2022. To elaborate a bit, although it should be obvious: OR formulations involve genuine physical non-unitarity in some form, such that the measurement process is defined at the appropriate micro-level and an interaction constituting a ‘measurement’ according to that theory results in an objectively existing outcome such that the measured system is accurately described by the corresponding eigenstate. This means that an OR formulation (such as the transactional interpretation, TI and its relativistic version RTI; cf. Kastner 2022) would dictate that Schrödinger’s Cat is not in a macroscopic superposition, since there definitely is or is not an atomic decay in the box at some particular time (cf. Kastner and Cramer 2018). It would similarly dictate that Bob and his lab are not in a macroscopic superposition under the envisioned circumstances of Bob’s measurement. Thus Wigner, under an OR theory, will accordingly predict that his friend Bob will actually get an outcome corresponding to an eigenstate for his measured system, and Wigner will assign the appropriate proper mixed state to Bob’s system if Wigner does not currently know that outcome. Wigner will not describe Bob and his lab by a macroscopic superposition, for the OR theory would not predict that (as discussed above). Only CQM predicts that. As noted above, whether or not one happens to like or accept or believe a particular OR theory to be fully adequate is irrelevant here: such a theory specifically says that measurements involve collapse and predicts that superpositions therefore do not persist beyond that point. And indeed Adlam acknowledges in (2023) that one such OR theory (the Transactional Interpretation)“solves the measurement problem,” as noted above. So one cannot purport to be applying an OR theory while assuming the existence of macroscopic superpositions denied by OR theories. Yet, as we will see in what follows, AR2022 do just that.
@@ 2. Critique
The above quote from AR2022, namely: “Unitary quantum mechanics does not provide any mechanism for a single measurement outcome to be selected and actualised for Alice in the first place….” illustrates the conventional assumption of CQM that the actual physical content of quantum theory must be restricted to unitary evolution, and therefore that measurement outcomes are not accounted for by the theory but are brute matters of fact for which some ‘interpretational’ account is needed. However, before one even gets to issues of consistency or correspondence, it is worth noting that RQM simply helps itself to the observed fact, admitted by the authors as not at all accounted for by CQM, that measurements have outcomes—even if only ‘relative’ to some observer or another. This is not a matter of the theory being ‘interpreted’ but merely an ad hoc patching of its key, acknowledged lacuna, since the theory provides for no outcomes, ‘relative’ or otherwise. In this sense, RQM is not really an interpretation of CQM but a different theory consisting of CQM along with an ad hoc postulate in place of the projection postulate; namely “outcomes occur relative to observers.” The latter is not a physical postulate that serves to explain empirical data, but rather simply postulates the empirical data that its underlying theory admittedly fails to explain. RQM thus postulates the explanandum, and that makes RQM vacuous as an explanans.
In contrast, OR theories, as noted above, provide for the actualization of an outcome under specific physical circumstances involving some form of non-unitarity, and are not observer-dependent; thus in OR theories there is never an ‘inner’ or ‘outer’ observer’ as in the WF scenario. There is a well-defined point at which collapse/reduction occurs as an invariant physical event, and any observer describes the situation accordingly. However, CQM is subject to the WF and related experiments, in which an ‘inner observer’ and ‘outer observer’ disagree on the states of the systems involved, which raises the possibility, made explicit in the Frauchiger-Renner paradox, that the ‘inner observer’ could be found to make a ‘wrong’ prediction. The latter circumstance refutes the central postulate of RQM, stated in Laudisa and Rovelli (2021) as follows:
@@
The form of RQM reviewed and slightly revised in AR2022—call it RQM22—thus abandons the central postulate of RQM (although the present author has been unable to find any admission on the part of RQM advocates of the fact that its core assumption is violated by the CQM that they purport to be interpreting). The original theoretically contradicted postulate is now replaced by the following five postulates:
@@
5. Internally consistent descriptions: In a scenario where F measures S, and W also measures S in the same basis, and W then interacts with F to ‘check the reading’ of a pointer variable (i.e. by measuring F in the appropriate ‘pointer basis’), the two values found are in agreement. (quoted from AR2022).
@@
These postulates introduce undefined terms such as the ‘fact’ and ‘event’ of Postulate 1. As mentioned above, if these are supposed to denote measurement outcomes, the authors help themselves to the occurrence of outcomes ‘for observers’ despite the fact, acknowledged by them, that “Unitary quantum mechanics does not provide any mechanism for a single measurement outcome to be selected and actualised for Alice in the first place…”. While AR2022 go on to talk about “flashes” as a matter of purported ontology, under CQM these ad hoc entities can never constitute the “observer-independent events” offered, as we will see in what follows. In addition, Postulate 3’s purpose seems unclear. If the intent is to rule out an observer-independent account, one might wonder why that would be necessary in view of the authors’ claim that their interpretation yields observer-independent facts. [^5] Deeming certain questions “meaningless” can serve to evade legitimate questions; in particular, comparison of accounts would seem to be meaningful if one is purporting to provide a physical and even ‘observer-independent’ ontology of the world, which AR2022 are doing.
@@
As an aside, the present author certainly agrees with the authors that one should aim for a fully relativistic interpretation, one which takes into account that spacetime could be emergent from the quantum level and thus is not assumed as a fixed background or container for events, as in their remark: “there is an alternative view which suggests that in fact, spacetime should be understood to emerge from a background of quantum events.” Indeed a comprehensive account of just how this could work is provided in Schlatter and Kastner (2023). The quote from Born is also apt: “One should not transfer the concept of space-time as a four-dimensional continuum from the macroscopic world of common experience to the atomistic world; manifestly the latter requires a different type of manifold.” However, these considerations do not save RQM, nor its recent reworking in AR2022, from the problems discussed herein.
@@
The authors review the Wigner’s Friend scenario to which CQM is subject, in which an ‘inner observer’ (Bob) performs a measurement M, finds an outcome and attributes the relevant eigenstate to his system, while Wigner (the ‘outer observer’) describes Bob with a coherent superposition, so that the two disagree on what quantum theory predicts. They claim that this disagreement cannot be corroborated in any way: “Wigner and Bob simply live within incommensurate realities now, and no attempt to reach across and bridge the gap can possibly succeed.”. (AR2022)
@@
“Definition 4.1. Cross-perspective links: In a scenario where some observer Alice measures a variable V of a system S, then provided that Alice does not undergo any interactions which destroy the information about V stored in Alice’s physical variables, if Bob subsequently measures the physical variable representing Alice’s information about the variable V, then Bob’s measurement result will match Alice’s measurement result.” (AR20220)
However, 4.1 does not succeed in remedying the essential inconsistency faced by CQM, as we will see in what follows. We now elucidate four consequential errors made by the authors:
@@
This is a very peculiar statement. The assertion “Wigner is wrong” arises only from the authors’ refusal to allow Wigner to apply the very OR formulation they claim to be applying. As discussed above, OR instructs Wigner not to describe Bob by a macroscopic superposition but by a proper mixed state. The authors even admit this: “those [OR] interpretations affirm that there will be an actual or effective collapse of the wavefunction at around the time of Bob’s measurement so that Bob will no longer be in a coherent superposition by the time of Wigner’s measurement.” So why do they assume that Wigner will describe Bob by a coherent superposition dictated not by OR but by CQM? Perhaps more to the point, why do the authors ignore the fact, just stated by them, that the OR theory dictates that “Bob will no longer be in a coherent superposition by the time of Wigner’s measurement”? The fact that OR theories preclude the WF-type paradoxes is acknowledged by Baumann and Brukner (2019, 3), who say:
“[OR] theories postulate a modification of standard quantum mechanics, such as spontaneous [7] and gravity-induced [4, 8] collapse models, which become significant at the macroscopic scale. These models could exclude superpositions of observers as required in Wigner’s friend-type experiments and hence the protocol cannot be run.”  [^6] Baumann and Brukner go on to say “In our view this is the most radical position as it would give rise to new physics..”. However, the widespread assumption that OR theories necessarily depart from the predictions of quantum theory is incorrect, since at least one OR-type theory, the Transactional Interpretation, not only does not depart from the Born Rule but physically derives it. So any “new physics” of TI is confined to the very old but overlooked physics of the direct-action theory of fields, which yields an empirically equivalent form of quantum theory that also accounts for “measurement” (cf. Kastner 2022, Chapter 5).
Now, if one thought a particular OR formulation failed to accomplish actual reduction at a macroscopic level, one could quibble with that. But that is not what AR2022 have done. They admit that an OR theory precludes macroscopic superpositions. Yet they unaccountably persist in assuming that the WF scenario still applies, falsely implying that the OR theory does not give a satisfying resolution of a ‘paradox’ to which it is not even subject, while their proposed interpretation supposedly does.
@@
The inconsistencies that arise in the MP-generated paradoxes involve supposedly observed facts or predictions that are in conflict among different observers. Many discussions of these inconsistencies treat them as “incommensurable realities.” I will term such claims IRC, the “incommensurate realities” claim, exemplified by AR2022 as above:
@@
(IRC) “Wigner and Bob simply live within incommensurate realities now, and no attempt to reach across and bridge the gap can possibly succeed.”
AR then attempt to remedy that asserted situation with what they term ‘cross-perspective links’ (CPL) as proposed in Postulate 4.1. However, this revision does nothing to remedy the fact that the ‘inner observer’ can indeed find that his prediction deviates from that of the ‘outer observer’ (according to CQM). This is demonstrated in an extant counterexample to IRC (discussed in Kastner 2020, 2021, 2022, and Baumann and Brukner 2019). The following material is excerpted from Kastner 2022. The term ‘UOA’ stands for ‘unitary outcome assumption,’ which is the standard working assumption under CQM that outcomes (somehow) result from a unitary interaction alone (even though, as noted above, CQM fails to account for outcomes).
{addmargin}
[2em]2em
@@
The common assumption that inconsistencies involving specific outcomes must always remain hidden, based on the idea that they are only observer-relative, is untenable via a specific counterexample. The example (similar to a scenario of Baumann and Brukner, 2019) demonstrates that attributing outcomes to entangled subsystems, even if designated as only “relative” to the subsystem or involving only “subjective collapse,” (as discussed by Baumann and Wolf, 2018), leads to empirical inconsistency. [^7] Deutsch (1985) considers a scenario with some similarities, but presents it as an empirical comparison between the Everettian picture and the Copenhagen interpretation. The present work considers implications of this type of empirical disagreement that do not seem to have yet been fully appreciated.
Assume, as in Proietti et al (2019), that microscopic quantum systems such as atoms, molecules, and photons can act as “observers” whose correlations with other similar degrees of freedom constitute “measurements” (this is the UOA). Let $W$ and $F$ comprise degrees of freedom subject to 2D state spaces, such as spin-$\frac{1}{2}$ atoms. Let us suppose that $F$ is “measuring” another spin-$\frac{1}{2}$ system $A$, prepared in equal superposition of outcomes ‘up’ and ‘down’ along a direction $Z$, via a Stern-Gerlach device. $F$ comprises 2 degrees of freedom: $B$, acting as a pointer/memory, and $C$, for communication. $C$’s possible states are ‘ground’ and ‘excited,’ where the latter would be triggered by a photon signal from $W$. While this is a thought experiment and does not depend on any particular experimental setup, one could think of $F$ as a micro-detector whose $B$ states are correlated with $A$ via an appropriate interaction Hamiltonian. $C$ remains in its initial unexcited state $|0\rangle$ at this stage. Thus, based on the idea that an outcome is relative to an observer, according to $F$ after the “measurement” of $A$ by $B$, $A$ the relevant degrees of freedom are either in the state:

### v2

@@
ABSTRACT. Quantum theory in its conventional formulation is notoriously subject to various measurement-related paradoxes, as exemplified by the “Schrödinger’s Cat” and “Wigner’s Friend” thought experiments. It has been shown, for example by Frauchiger and Renner, that nested measurements such as those occurring in the Wigner’s Friend experiment can lead to inconsistencies concerning the putative outcomes of measurements. Such inconsistencies are commonly presumed to remain private and incommensurable, but this is not the case. A counterexample, in which the inconsistencies can be revealed among the observers, is reviewed. The implications for a recent attempt to shield Relational Quantum Mechanics from such inconsistencies are considered, and it is concluded that the attempt is not successful. Further implications for the state of the debate concerning the viability of quantum theory in its various formulations are discussed.
@@ 1. Introduction and Background
This paper reviews the challenge to conventional quantum theory from inconsistencies arising from its prediction of macroscopic superpositions such as Schrödinger’s Cat and Wigner’s Friend. I focus on an interpretational approach first proposed by Rovelli (1996) called “Relational Quantum Mechanics”(henceforth RQM). RQM is revisited with modifications in Adlam and Rovelli 2022. The latter paper is the subject of the present critique.
First, let us clarify what is meant by “conventional quantum theory” for purposes of the present work. This characterization is meant to encompass all formulations of quantum theory that provide a quantitative, physical account only of unitarily evolving processes, i.e., those described by the Schrödinger equation. In view of that feature, we can for greater clarity use the term Unitary Quantum Theory to denote such formulations (which is consistent with terminology in Adlam and Rovelli 2022). This term applies not only to Everettian approaches (which deny non-unitary collapse or reduction), but also to so-called “textbook quantum theory.” Although the latter makes use of a projection postulate involving non-unitary collapse, that is traditionally invoked as a purely calculational step for empirical adequacy and utility. The postulate provides no quantitative physical account for a “measurement” instantiating non-unitary evolution. Thus, quantum theory’s physical (as opposed to mathematical/formal) content is taken as limited to unitary evolution in both Everettian and textbook formulations. Under this broad class of formulations, one can be agnostic about whether the projection postulate should be part of quantum theory, but one assumes that the theory “really” has only unitary evolution in a physically quantifiable sense. And that is the prevailing convention.
The projection postulate was termed Process 1 by Von Neumann (1932), and consists of two distinct mathematical stages. The first is the transition of a pure quantum state, represented by a Hilbert space vector, to a proper mixed state, represented by a density operator on Hilbert space. The second stage is the reduction or collapse to a pure state eigenvector of the measured observable. However, as noted above, in textbook quantum theory, Process 1 has no quantitative physical account that specifies how such a transition, lying outside the Schrödinger evolution, occurs. (The unitary Schrödinger evolution was called Process 2 by Von Neumann.)
The textbook theory leads to the so-called “Shifty Split” in which Schrödinger unitary evolution is applied to some subset of all systems in the universe—the system of interest—while everything else is taken as outside the domain of applicability of quantum theory (showing again that the theory is assumed to be unitary-only in actual physical content). This places the “observer” outside the theory, yet the theory itself cannot specify what constitutes an “observer” as a physical system. Thus the split becomes an ad hoc device that is applied without a well-defined physical justification. This situation is an aspect of the measurement problem, i.e., the fact that the theory cannot quantitatively account for the occurrence of outcomes of measurements describable by eigenstates of the measured observable, since the latter are in general not arrived at through unitary evolution. [^1] Everettian approaches often claim to resolve the measurement problem, but they depend strongly on decoherence arguments which ultimately involve taking improper mixtures as providing justification for outcomes, an arguably illicit step. This approach has been critiqued in Kastner (2014), which raises a circularity objection, and Kastner (2022), which raises an inconsistency objection.
@@
The measurement problem of Unitary Quantum Theory is notoriously exemplified by such paradoxical thought experiments as Schrödinger’s Cat and Wigner’s Friend, which involve “absurd” macroscopic superpositions (the characterization in Frauchiger and Renner 2018, and clearly Schrödinger’s intent). Since the theory (even if utilizing a “projection postulate”) denies any real physical non-unitarity, it must treat observed determinate outcomes as brute empirical facts with no physical explanation within the theory. More strongly, observed outcomes fail to be predicted by the theory, as the macroscopic superposition paradoxes illustrate. This point is granted by Adlam and Rovelli 2022, who say (in the context of a Wigner’s Friend experiment):
@@
“Unitary quantum mechanics does not provide any mechanism for a single measurement outcome to be selected and actualised for Alice in the first place, so it certainly cannot tell us anything about the relationship between her outcome and Bob’s outcome - this is a question which lies entirely outside the unitary part of the theory…” (Adlam and Rovelli 2022)
@@
The locution “unitary part of the theory” might seem to suggest that there is another “part of the theory,” but, as discussed above, that merely leaves room for an ad hoc application of a collapse postulate, or other means of stipulating an outcome, which is subject to the Shifty Split between the system to which quantum theory is taken as applying and some putative “outside observer.” This situation is what gives rise to the Wigner’s Friend and related inconsistencies such as that of the Frauchiger-Renner Paradox (Frauchiger and Renner, 2018). Of course, there are many attempts to deal with the measurement problem and resulting inconsistencies arising from Unitary Quantum Theory. Adlam and Rovelli 2022 is one of those attempts; I argue in what follows that it fails.
@@
In contrast to Unitary Quantum Theory, formulations involving objective reduction or collapse (henceforth OR) are not subject to the measurement problem and thus do not lead to macroscopic superpositions such as that of Bob and his lab in the Wigner’s Friend paradox. [^2] Some examples are the gravitational collapse model of Penrose 2000, the Ghirardi-Rimini-Weber (1986) model, and the Transactional Interpretation (Cramer 1986, most recent relativistic version in Kastner 2022). The latter is distinct from the two former models in that it does not modify the Schrödinger equation but uses a different theory of fields that naturally leads to a non-unitary dynamics. This is the case regardless of one’s opinion of the theoretical or practical adequacy of such theories, which is a separate issue. Interestingly, one of the authors of Adlam and Rovelli 2022 (EA) does acknowledge that a particular OR formulation solves the measurement problem in a footnote to a paper entitled “Do We Have Any Viable Solution to the Measurement Problem?”, where she says:
@@
“…the transactional interpretation…is a solution to the measurement problem which has the property of being retrocausal.” (Adlam 2023). [^3] Adlam’s references are out of date, however; in addition to Cramer (1986), a more recent, fully relativistic account is Kastner (2022); see also Kastner (2018) in regard to TI’s solution to the measurement problem.
@@
(This comment was made as an aside in the context of properties a particular formulation or interpretation might have. Oddly, however, although acknowledged by the author as a solution to the measurement problem in the quoted footnote, the transactional interpretation is not mentioned anywhere in the body of the paper, despite its title—even with the emphasis on “Any”).
@@
Thus, at the risk of repetition (for some emphasis seems necessary), OR formulations eliminate the measurement problem so that such paradoxes as Schrödinger’s Cat and Wigner’s Friend do not arise in the first place. It is important to keep this point—that the notorious measurement-related paradoxes do not arise under OR formulations—firmly in mind, since it appears to have been somehow forgotten in Adlam and Rovelli 2022. To elaborate a bit, although it should be obvious: OR formulations involve genuine physical non-unitarity in some form, such that the measurement process is defined at the appropriate micro-level and an interaction constituting a ‘measurement’ according to that theory results in an objectively existing outcome such that the measured system is accurately described by the corresponding eigenstate. This means that an OR formulation (such as the transactional interpretation, TI and its relativistic version RTI; cf. Kastner 2022) would dictate that Schrödinger’s Cat is not in a macroscopic superposition, since there definitely is or is not an atomic decay in the box at some particular time (cf. Kastner and Cramer 2018). It would similarly dictate that Bob and his lab are not in a macroscopic superposition under the envisioned circumstances of Bob’s measurement. Thus Wigner, under an OR theory, will accordingly predict that his friend Bob will actually get an outcome corresponding to an eigenstate for his measured system, and Wigner will assign the appropriate proper mixed state to Bob’s system if Wigner does not currently know that outcome. Wigner will not describe Bob and his lab by a macroscopic superposition, for the OR theory would not predict that (as discussed above). Only Unitary Quantum Theory predicts that. As noted above, whether or not one happens to like or accept or believe a particular OR theory to be fully adequate is irrelevant here: such a theory specifically says that measurements involve collapse and predicts that superpositions therefore do not persist beyond that point. And indeed Adlam acknowledges in (2023) that one such OR theory (the Transactional Interpretation)“solves the measurement problem,” as noted above. So one cannot purport to be applying an OR theory while assuming the existence of macroscopic superpositions precluded by OR theories. Yet, as we will see in what follows, Adlam and Rovelli 2022 do just that.
@@ 2. Critique
The above quote from Adlam and Rovelli 2022, namely: “Unitary quantum mechanics does not provide any mechanism for a single measurement outcome to be selected and actualised for Alice in the first place….” illustrates the conventional assumption of Unitary Quantum Theory that the actual physical content of quantum theory must be restricted to unitary evolution, and therefore that measurement outcomes are not accounted for by the theory but are brute matters of fact for which some ‘interpretational’ account is needed. However, before one even gets to issues of consistency or coherence, it is worth noting that RQM simply helps itself to the observed fact, admitted by the authors as not at all accounted for by Unitary Quantum Theory, that measurements have outcomes—even if only ‘relative’ to some observer or another. This is not a matter of the theory being ‘interpreted’ but rather an ad hoc patching of its key, acknowledged lacuna, since the presumed unitary-only physical evolution of the conventional theory itself licenses no outcomes, ‘relative’ or otherwise. In this sense, RQM simply replaces the usual ad hoc projection postulate with a slightly different ad hoc quasi-projection postulate: namely, “outcomes occur relative to observers.” It is important to recognize that the latter is not a physical postulate that serves to explain empirical data, but instead simply appropriates by fiat the very empirical data that its underlying theory admittedly fails to explain. RQM thus postulates the explanandum, and that makes RQM vacuous as an explanans.
In contrast, OR theories, as noted above, provide for the actualization of an outcome under specific physical circumstances involving some form of non-unitarity, and are not observer-dependent. Thus, there is a well-defined point at which collapse/reduction occurs, yielding an outcome as an invariant physical event, and any observer describes the situation accordingly. This precludes the Wigner’s Friend and related “nested observer” scenarios. However, Unitary Quantum Theory remains subject to such scenarios, which raise the possibility, made explicit in the Frauchiger-Renner paradox, that the ‘inner observer’ could be found to make a ‘wrong’ prediction. The latter circumstance refutes the central postulate of RQM, stated in Laudisa and Rovelli (2021) as follows:
@@
The form of RQM reviewed and slightly revised in Adlam and Rovelli 2022—call it RQM22—thus abandons the central postulate of RQM (although the present author has been unable to find any admission on the part of RQM advocates of the fact that its core assumption is violated by the Unitary Quantum Theory that they purport to be interpreting). The original theoretically contradicted postulate is now replaced by the following five postulates:
@@
5. Internally consistent descriptions: In a scenario where F measures S, and W also measures S in the same basis, and W then interacts with F to ‘check the reading’ of a pointer variable (i.e. by measuring F in the appropriate ‘pointer basis’), the two values found are in agreement. (quoted from Adlam and Rovelli 2022).
@@
These postulates introduce undefined terms such as the ‘fact’ and ‘event’ of Postulate 1. As mentioned above, if these are supposed to denote measurement outcomes, the authors help themselves to the occurrence of outcomes ‘for observers’ despite the fact, acknowledged by them, that “Unitary quantum mechanics does not provide any mechanism for a single measurement outcome to be selected and actualised for Alice in the first place…”. [^5] The fact that RQM helps itself to the notion of “measurement” and “measurement outcomes” not licensed by Unitary Quantum Theory is demonstrated in Rovelli (1998), which dodges the issue entirely by saying “For some reason, at some point we have to (or we can) replace the pure state $\Psi(T)$ with a mixed state.” It is then simply stipulated without justification that the system acquires some eigenvalue of the observable in question. While the authors go on to talk about “flashes” as a matter of purported ontology, under Unitary Quantum Theory these ad hoc entities can never constitute the “observer-independent events” offered, as we will see in what follows. In addition, Postulate 3’s purpose seems unclear. If the intent is to rule out an observer-independent account, one might wonder why that would be necessary in view of the authors’ claim that their interpretation yields observer-independent facts. [^6] Deeming certain questions “meaningless” can serve to evade legitimate questions; in particular, comparison of accounts would seem to be meaningful if one is purporting to provide a physical and even ‘observer-independent’ ontology of the world, which Adlam and Rovelli are doing.
@@
As an aside, the present author certainly agrees with the authors that one should aim for a fully relativistic interpretation, one which takes into account that spacetime could be emergent from the quantum level and thus is not assumed as a fixed background or container for events, as in their remark: “there is an alternative view which suggests that in fact, spacetime should be understood to emerge from a background of quantum events.” Indeed a comprehensive account of just how this could work is provided in Schlatter and Kastner (2023). The quote from Born is also apt: “One should not transfer the concept of space-time as a four-dimensional continuum from the macroscopic world of common experience to the atomistic world; manifestly the latter requires a different type of manifold.” However, these considerations do not save RQM, nor its recent reworking in Adlam and Rovelli 2022, from the problems discussed herein.
@@
The authors review the Wigner’s Friend scenario to which Unitary Quantum Theory is subject, in which an ‘inner observer’ (Bob) performs a measurement M, finds an outcome and attributes the relevant eigenstate to his system, while Wigner (the ‘outer observer’) describes Bob with a coherent superposition, so that the two disagree on what quantum theory predicts. They claim that this disagreement cannot be corroborated in any way: “Wigner and Bob simply live within incommensurate realities now, and no attempt to reach across and bridge the gap can possibly succeed.”. (Adlam and Rovelli 2022)
@@
“Definition 4.1. Cross-perspective links: In a scenario where some observer Alice measures a variable V of a system S, then provided that Alice does not undergo any interactions which destroy the information about V stored in Alice’s physical variables, if Bob subsequently measures the physical variable representing Alice’s information about the variable V, then Bob’s measurement result will match Alice’s measurement result.” (Adlam and Rovelli 2022)
However, 4.1 does not succeed in remedying the essential inconsistency faced by Unitary Quantum Theory, as we will see in what follows. We now elucidate four consequential errors made by the authors:
@@
This is a very peculiar statement. The assertion “Wigner is wrong” arises only from the authors’ refusal to allow Wigner to apply the very OR formulation they claim to be applying. As discussed above, OR instructs Wigner not to describe Bob by a macroscopic superposition but by a proper mixed state. The authors even admit this: “those [OR] interpretations affirm that there will be an actual or effective collapse of the wavefunction at around the time of Bob’s measurement so that Bob will no longer be in a coherent superposition by the time of Wigner’s measurement.” So why do they assume that Wigner will describe Bob by a coherent superposition dictated not by OR but by Unitary Quantum Theory? Perhaps more to the point, why do the authors ignore the fact, just stated by them, that the OR theory dictates that “Bob will no longer be in a coherent superposition by the time of Wigner’s measurement”?
The fact that OR theories preclude the Wigner’s Friend-type paradoxes is acknowledged by Baumann and Brukner (2019, 3), who say:
“[OR] theories postulate a modification of standard quantum mechanics, such as spontaneous [7] and gravity-induced [4, 8] collapse models, which become significant at the macroscopic scale. These models could exclude superpositions of observers as required in Wigner’s friend-type experiments and hence the protocol cannot be run.”  [^7] Baumann and Brukner go on to say “In our view this is the most radical position as it would give rise to new physics.” However, the widespread assumption that OR theories necessarily depart from the predictions of quantum theory is incorrect, since at least one OR-type theory, the Transactional Interpretation, not only does not depart from the Born Rule but physically derives it. So any “new physics” of TI is confined to the very old but overlooked physics of the direct-action theory of fields, which yields an empirically equivalent form of quantum theory that also accounts for “measurement” (cf. Kastner 2022, Chapter 5).
Now, if one thought a particular OR formulation failed to accomplish actual reduction at a macroscopic level, one could quibble with that. But that is not what the authors have done. They admit that an OR theory precludes macroscopic superpositions. Yet they unaccountably persist in assuming that the Wigner’s Friend scenario still applies, falsely implying that the OR theory does not give a satisfying resolution of a ‘paradox’ to which it is not even subject, while their proposed interpretation supposedly does.
@@
The inconsistencies that arise in the MP-generated paradoxes involve supposedly observed facts or predictions that are in conflict among different observers. Many discussions of these inconsistencies assume that these putative observed facts cannot be compared among observers of different “nesting” levels, and thus constitute “incommensurate realities.” This assumption exemplified in Adlam and Rovelli 2022 in their comment:
@@
“Wigner and Bob simply live within incommensurate realities now, and no attempt to reach across and bridge the gap can possibly succeed.”
@@
The authors then attempt to remedy that asserted situation with what they term ‘cross-perspective links’ as proposed in Postulate 4.1. However, this revision does nothing to remedy the fact that the ‘inner observer’ can indeed find that his prediction deviates from that of the ‘outer observer’ (according to Unitary Quantum Theory). This is demonstrated in an extant example discussed in Kastner 2020, 2021, 2022, and Baumann and Brukner 2019, which we now review (closely following the account in Kastner, 2022). The example, a variant of the Wigner’s Friend experiment, demonstrates that attributing outcomes to entangled subsystems, even if designated as only “relative” to the subsystem or involving only “subjective collapse,” (as discussed by Baumann and Wolf, 2018), leads to overt empirical inconsistency as opposed to “hidden” and “incommensurate” inconsistencies as commonly assumed. [^8] Deutsch (1985) considers a similar scenario, but offers it as an empirical comparison between the Everettian picture and the Copenhagen interpretation. Here I discuss implications of this type of empirical disagreement that do not yet seem to have been fully appreciated.
First, recall that Unitary Quantum Theory has no well-defined criterion for what sort of system counts as a “measurement device” to which an outcome could be attributed, nor for what sort of system counts as an “observer” who could be taken as acquiring knowledge of an outcome. This leads to the common treatment, as for example in Proietti et al (2019), in which a microscopic quantum system such as an atom, molecule, or photon is taken as an “observer” whose correlation with another degree of freedom can (when desired) constitute a “measurement” yielding an outcome. With that background, consider a Wigner’s Friend experiment in which the two parties, labeled $W$ and $F$, are instantiated by degrees of freedom subject to 2D state spaces, such as spin-$\frac{1}{2}$ atoms. As usual, $F$ is taken as an “inner observer” and $W$ as an “outer observer.” $F$ comprises 2 degrees of freedom: $B$, acting as a pointer/memory, and $C$, for communication. $C$’s possible states are ‘ground’ and ‘excited,’ where the latter would be triggered by a photon signal from the outer observer $W$.
Let us suppose that $F$ is measuring another spin-$\frac{1}{2}$ system $A$, prepared in an equal superposition of states ‘up’ and ‘down’ along a direction $Z$, via a Stern-Gerlach device. One can think of $F$ as a micro-detector whose $B$ states become correlated with $A$ via an appropriate interaction Hamiltonian. $C$ remains in its initial unexcited state $|0\rangle$ at this stage.
Now, based on the conventional approach of attributing an outcome to a system based on the establishment of a suitable interaction Hamiltonian as above, according to $F$ after the “measurement” of $A$ by $B$, $A$ the relevant degrees of freedom are either in the state: